# adversarial attack DB Diffusion Buffer

### v1

@@ 1 Introduction
While adversarial attacks are most commonly designed for classification systems such as ASR, they can also be applied to regression tasks. For instance, one could think of a scenario where adversarial noise is added by an attacker, such that a Speech Enhancement (SE) system of a speech communication device such as a hearing aid or a telephony system outputs speech with a completely different and unintended semantic meaning. However, modifying a speech signal such that the semantic meaning is deliberately changed by a SE system has received considerably less attention, also because traditional speech enhancement systems, e.g. based on Wiener filtering, are hardly expressive enough to change the semantic meaning of speech. However, now that both predictive and in generative SE systems are becoming increasingly expressive, the risk of adversarial attacks increases. It is therefore important to investigate such potential vulnerabilities in more detail.
Therefore, in this paper, we explore the possibility of targeted adversarial attacks on modern expressive predictive and generative SE systems to produce an output with a completely different meaning than the original speech signal had. We consider a white-box scenario in which the adversary has full knowledge of the enhancement model and can compute gradients through it. We use a psychoacoustic model to make the adversarial noise as inaudible as possible, while keeping the intelligibility and quality of the enhanced speech as high as possible. We draw inspiration from psychoacoustic masking attacks in ASR[3], extending them to the SE domain and addressing new challenges that arise, especially for diffusion-based generative models. The key contributions of this work are as follows: (1) We propose an attack loss function and optimization procedure applicable to both traditional and diffusion-based SE, incorporating psychoacoustic constraints. (2) We analyze the differences in vulnerability between a standard predictive models like a regression SE, a mask-based SE, and a score-based diffusion SE (SGMSE+), including the effect of stochastic sampling in diffusion. (3) We evaluate on EARS-WHAM-v2 using two metric groups: attack success: DistillMOS and Word Error Rate (WER), Extended Short-Term Objective Intelligibility (ESTOI), Perceptual Objective Listening Quality Analysis (POLQA) against the target clean speech) and perturbation impact POLQA, ESTOI between the attacked and original mixtures, plus the perturbation signal-to-noise ratio (SNR).
@@ 2 Background
We introduce an SE model, denoted as $f_{SE}$, which serves as a generic operator that takes a noisy STFT $Y$ as input and aims to estimate the clean speech signal $S$, producing an output $\hat{S}_{Y}$.
$\hat{S}_{Y}=f_{SE}(Y)$ (1)
@@ 3 Adversarial Attacks
$\hat{S}(Y^{\mathrm{user}}+\delta)\;=\;f_{SE}(Y^{\mathrm{user}}+\delta\big).$ (3)
Then we optimize $\delta$ to such that the SE system’s output is close to $S^{\mathrm{attacker}}$ instead of $S^{\mathrm{user}}$using an MSE loss function
$\mathcal{L}_{\mathrm{adv}}(\delta)\;=\;\sum_{q,n}\big|\hat{S}_{q,n}(Y^{\mathrm{user}}+\delta)-S^{\mathrm{attacker}}_{q,n}\big|^{2}.$ (4)
@@ 3.2 Psychoacoustic Hiding
$H_{q,n}$ is the per-time-frequency audibility limit: it combines the absolute threshold in quiet with signal-dependent masking, so that perturbations kept below $H$ are hidden for the listener. Following [3], we estimate $H$ from noisy observation $Y^{\mathrm{user}}$ with the MPEG-1 psychoacoustic model. For reproducibility, we use an open source implementation [^2] https://github.com/RUB-SysSec/dompteur/tree/main/standalone-psychoacoustic-filtering.
@@
This yiels a per-bin deviation in dB and we can directly compare it to $H_{q,n}$
@@ 3.3 Projected Gradient Descent with an $\ell_{2}$ budget
$\Pi_{B_{2}(\varepsilon)}(\delta)=\delta\cdot\min\!\Big\{1,\;\frac{\varepsilon}{\|\delta\|_{2}}\Big\},\qquad\|\delta\|_{2}^{2}=\sum_{q,n}\big|\delta_{q,n}\big|^{2}.$
PGD solves the constrained attack by alternating a gradient step with a projection back onto the norm ball, such that in every iteratation $\|\delta\|_{2}\leq\varepsilon$ is satisfied. As a consequence, PGD limits the total energy of the attack, while $\hat{\Phi}$ spectrally shapes the attack such that it is masked by $Y^{\mathrm{user}}$.
@@
$\hat{S}$ $\hat{S}$ vs. $S^{\mathrm{attacker}}$ ($Y^{\mathrm{user}}{+}\delta$) vs. $Y^{\mathrm{user}}$ $Y^{\mathrm{user}}$$\delta$
@@
Table 1: Targeted-attack results on 100 EARS-WHAM-v2 pairs. Left block of metrics reports attack success for the enhanced signal $\hat{S}$ against the target $S^{\mathrm{attacker}}$; right block reports perturbation impact for $(Y^{\mathrm{user}}{+}\delta)$ against $Y^{\mathrm{user}}$ and the SNR of $\delta$ w.r.t. $Y^{\mathrm{user}}$ (higher SNR = quieter perturbation).
@@ 4.1 Model preparation
We took the SGMSE++ pipeline [^3] https://github.com/sp-uhh/sgmse as our baseline repository to obtain a strong diffusion-based SE system. We prepared a generative diffusion model and two predictive baselines that reuse the same backbone but differ only in how $\hat{S}$ is produced (Section 2). We took the NCSN++ U-Net backbone as a baseline architecture [9]. To reduce the computational complexity, we modified the network depth by decreasing the number of residual blocks from 2 to 1. Unless noted otherwise, the remaining SGMSE++ settings are left unchanged.
@@
  • Direct Map - a predictive network $d_{\theta}$ that maps $Y$ to $\hat{S}$, trained with the pointwise complex MSE to the clean target $S$.
@@ 4.2 Adversarial Attack with Psychoacoustic Hiding
From the EARS-WHAM-v2 test set (850 utterances; 6 speakers, 3 male / 3 female) we sample $100$ random pairs $(Y^{\mathrm{user}},S^{\mathrm{attacker}})$ of comparable duration, where $Y^{\mathrm{user}}$ is a noisy mixture from utterance 1 and $S^{\mathrm{attacker}}$ is a clean utterance 2. For each pair, we optimize a complex-valued perturbation $\delta\in\mathbb{C}^{F\times T}$ and construct the attacked input $Y^{\mathrm{user}}+\delta$ and output $\hat{S}(Y^{\mathrm{user}}{+}\delta)$ (3). Starting with $\delta^{(0)}\!=\!0$, we update $\delta$ with the PGD rule (11), using K=150K{=}150 iterations of SGD with momentum $0.4$ and learning rate $0.1$, while keeping model parameters fixed.
For the diffusion-based SE, we set the number of reverse steps to N=25N{=}25. To disentangle stochasticity from optimization effects, we consider two inference modes during the attack: (1) Diffusion - the stochastic reverse SDE; (2) Diffusion-fix - the same reverse SDE with a fixed Brownian path across attack iterations.
@@ 4.3 Evaluation metrics
We evaluate adversarial attacks along two axes: attack success - how well the output signal $\hat{S}(Y^{\mathrm{user}}+\delta)$ matches the target speech $S^{\mathrm{attacker}}$; and perturbation impact - how audible the perturbation $\delta$ is within the original mixture $Y^{\mathrm{user}}$.
Attack success. We report WER, POLQA, ESTOI between $\hat{S}(Y^{\mathrm{user}}+\delta)$ and $S^{\mathrm{attacker}}$, and DistillMOS [13] for $\hat{S}(Y^{\mathrm{user}}+\delta)$. For WER, we transcribe with five Whisper [14] variants (Large, Large-v1, Large-v2, Large-v3, Large-v3-turbo) and average the resulting WERs for stability.
Perturbation impact. To quantify how much the input mixture is altered, we use POLQA and ESTOI as a distance metric between $(Y^{\mathrm{user}}+\delta)$ and $Y^{\mathrm{user}}$.
@@
Here $\|\cdot\|_{2}$ denotes the Euclidean norm. A higher SNR indicates that the adversarial attack if of lower power.
@@
Table 2: Ablation of diffusion SE on the same 100 pairs. Unless specified, the sampler uses N=25N{=}25 reverse steps and stochastic sampling; Fix noise denotes a frozen noise path. The baseline model is trained with $\sigma_{\max}{=}0.5$; rows that list $\sigma_{\max}$ are separately trained models.
@@ 5 Results and Analysis
Table 1 summarizes targeted attacks on 100 EARS–WHAM–v2 pairs. Below we highlight the main results.
@@
(1) Fixed energy ($\varepsilon{=}10$), sweep $\lambda$: with the energy budget fixed, SNR and perturbation–impact stay within a narrow band (SNR $\approx 8-13\ {dB}$; impact POLQA $\approx 2.55-3.19$; impact ESTOI $\approx 0.65-0.72$), while attack success shifts markedly—e.g., WER $\approx 0.20\to 1.37$, AS–ESTOI $\approx 0.77\to 0.22$, AS–POLQA $\approx 2.62\to 1.07$ (DistillMOS $\approx 3.27\to 1.72$) as $\lambda$ tightens ($40\to 0$).
(2) Fixed masking ($\lambda{=}20$), sweep $\varepsilon$: enlarging $\varepsilon$ improves alignment (WER $1.17\to 0.23$ from $\varepsilon{=}3\to\infty$) while increasing audibility (impact POLQA $4.59\to 2.40$; SNR $32.46\to 7.72$ dB). Practically, moderate settings (e.g., $\lambda{=}20$, $\varepsilon\in{10,15}$) sit in a balanced region (WER $\approx 0.20-0.17$, perturbation impact POLQA $\approx 3.19-2.26$, SNR $\approx 12.88-5.64$ dB).
CRM (predictive, mask). For identical $(\lambda,\varepsilon)$, perturbation impact matches Direct Map closely (e.g., at $\lambda{=}20,\varepsilon{=}10$: impact POLQA $3.14$ vs. $3.19$; SNR $12.90$ vs. $12.88$ dB). Attack success is consistently, but only slightly lower (e.g., POLQA $1.81$ vs. $1.57$, WER $0.23$ vs. $0.20$ at the same setting). The short conclusion is that masking approach with similar perturbation impact metrics produces a worse attack.
@@ 6 Conclusions
In this paper, we show that modern speech enhancement systems exhibit some vulnerability to adversarial attacks. That is, by adding adversarial noise, an attacker can trick a speech enhancement system to output a completely different signal than intended by the used. To that end, we build white-box attacks with psychoacoustic masking and an $\ell_{2}$ budget for the adversarial attack. We test predictive (direct map, complex mask) and diffusion SE on EARS-WHAM-v2 using success and audibility metrics.
We show that predictive SE is comparably easily attacked with a realtively a small $\ell_{2}$ budget for the adversarial noise. The complex mask variant behaves similarly to direct mapping but is somewhat less vulnerable. Diffusion SE is more resistant to targeted manipulation: even with a larger $\ell_{2}$ budget for the adversarial noise, the output aligns less well with the attacker’s target signal and remains more audible than predictive baselines. We show that stochastic sampling in the reverse diffusion process increases robustness to attacks even further.

### v2

@@ 1 Introduction
While adversarial attacks are most commonly designed for classification systems such as ASR, they can also be applied to regression tasks. For instance, one could think of a scenario where adversarial noise is added by an attacker, such that a Speech Enhancement (SE) system of a speech communication device such as a hearing aid or a telephony system outputs speech with a completely different and unintended semantic meaning. However, modifying a speech signal such that the semantic meaning is deliberately changed by a SE system has received considerably less attention, also because traditional speech enhancement systems [4], e.g. based on Wiener filtering, are hardly expressive enough to change the semantic meaning of speech. However, now that both predictive and generative SE systems are becoming increasingly expressive, the risk of adversarial attacks increases. It is therefore important to investigate such potential vulnerabilities in more detail.
Therefore, in this paper, we explore the possibility of targeted adversarial attacks on modern expressive predictive and generative SE systems to produce an output with a completely different meaning than the original speech signal had. We consider a white-box scenario in which the adversary has full knowledge of the enhancement model and can compute gradients through it. We use a psychoacoustic model to make the adversarial noise as inaudible as possible, while keeping the intelligibility and quality of the enhanced speech as high as possible. We draw inspiration from psychoacoustic masking attacks in ASR [3], extending them to the SE domain and addressing new challenges that arise, especially for diffusion-based generative models. The key contributions of this work are as follows: (1) We propose an attack loss function and optimization procedure applicable to both traditional and diffusion-based SE, incorporating psychoacoustic constraints. (2) We analyze the differences in vulnerability between a standard predictive models like a regression SE, a mask-based SE, and a score-based diffusion SE (SGMSE+), including the effect of stochastic sampling in diffusion. (3) We evaluate on EARS-WHAM-v2 using two metric groups: attack success: DistillMOS and Word Error Rate (WER), Extended Short-Term Objective Intelligibility (ESTOI), Perceptual Objective Listening Quality Analysis (POLQA) against the target speech; and perturbation impact: POLQA, ESTOI between attacked and original mixtures, plus perturbation signal-to-noise ratio (SNR).
@@ 2 Background
We introduce an SE model, denoted as $f_{SE}$, which serves as a generic operator that takes a noisy STFT $Y$ as input and aims to estimate the clean speech signal $S$, producing an output $\hat{S}^{Y}$.
$\hat{S}^{Y}=f_{SE}(Y)$ (1)
@@ 3 Adversarial Attacks
$\hat{S}^{Y^{\mathrm{user}}+\delta}\;=\;f_{SE}\big(Y^{\mathrm{user}}+\delta\big).$ (3)
Then we optimize $\delta$ so that the SE system’s output is close to $S^{\mathrm{attacker}}$ instead of $S^{\mathrm{user}}$using an MSE loss function
$\mathcal{L}_{\mathrm{adv}}(\delta)\;=\;\sum_{q,n}\big|\hat{S}^{Y^{\mathrm{user}}+\delta}_{q,n}-S^{\mathrm{attacker}}_{q,n}\big|^{2}.$ (4)
@@ 3.2 Psychoacoustic Hiding
$H_{q,n}$ is the per-time-frequency audibility limit: it combines the absolute threshold in quiet with signal-dependent masking, so that perturbations kept below $H$ are hidden for the listener. Following [3], we estimate $H$ from noisy observation $Y^{\mathrm{user}}$ with the MPEG-1 psychoacoustic model. For reproducibility, we use an open-source implementation [^2] https://github.com/RUB-SysSec/dompteur/tree/main/standalone-psychoacoustic-filtering.
@@
This yields a per-bin deviation in dB and we can directly compare it to $H_{q,n}$
@@ 3.3 Projected Gradient Descent with an $\ell_{2}$ budget
$\Pi_{B_{2}(\varepsilon)}(\delta)=\delta\cdot\min\!\Big\{1,\;\frac{\varepsilon}{\|\delta\|_{2}}\Big\},\qquad\|\delta\|_{2}=\sqrt{\sum_{q,n}\big|\delta_{q,n}\big|^{2}}.$
PGD solves the constrained attack by alternating a gradient step with a projection back onto the norm ball, such that in every iteration $\|\delta\|_{2}\leq\varepsilon$ is satisfied. As a consequence, PGD limits the total energy of the attack, while $\hat{\Phi}$ spectrally shapes the attack such that it is masked by $Y^{\mathrm{user}}$.
@@
$\hat{S}^{Y^{\mathrm{user}}+\delta}$ $\hat{S}^{Y^{\mathrm{user}}+\delta}$ vs. $S^{\mathrm{attacker}}$ ($Y^{\mathrm{user}}{+}\delta$) vs. $Y^{\mathrm{user}}$ $Y^{\mathrm{user}}$$\delta$
@@
Table 1: Targeted-attack results on 100 EARS-WHAM-v2 pairs. Left block of metrics reports attack success for the enhanced signal $\hat{S}^{Y^{\mathrm{user}}+\delta}$ against the target $S^{\mathrm{attacker}}$; right block reports perturbation impact for $(Y^{\mathrm{user}}{+}\delta)$ against $Y^{\mathrm{user}}$ and the SNR of $\delta$ w.r.t. $Y^{\mathrm{user}}$ (higher SNR = quieter perturbation).
@@ 4.1 Model preparation
We adopt the SGMSE++ pipeline [^3] https://github.com/sp-uhh/sgmse as our baseline repository to obtain a strong diffusion-based SE system. We prepared a generative diffusion model and two predictive baselines that reuse the same backbone but differ only in how $\hat{S}$ is produced (Section 2). We took the NCSN++ U-Net backbone as a baseline architecture [10]. To reduce the computational complexity, we modified the network depth by decreasing the number of residual blocks from 2 to 1. Unless noted otherwise, the remaining SGMSE++ settings are left unchanged.
@@
  • Direct Map - a predictive network $d_{\theta}$ that maps $Y$ to $\hat{S}$, trained with the point-wise complex MSE to the clean target $S$.
@@ 4.2 Adversarial Attack with Psychoacoustic Hiding
From the EARS-WHAM-v2 test set (850 utterances; 6 speakers, 3 male / 3 female) we sample $100$ random pairs $(Y^{\mathrm{user}},S^{\mathrm{attacker}})$ of comparable duration, where $Y^{\mathrm{user}}$ is a noisy mixture from utterance 1 and $S^{\mathrm{attacker}}$ is a clean utterance 2. For each pair, we optimize a complex-valued perturbation $\delta\in\mathbb{C}^{F\times T}$ and construct the attacked input $Y^{\mathrm{user}}+\delta$ and output $\hat{S}^{Y^{\mathrm{user}}+\delta}$ (3). Starting with $\delta^{(0)}\!=\!0$, we update $\delta$ with the PGD rule (11), using K=150K{=}150 iterations of SGD with momentum $0.4$ and learning rate $0.1$, while keeping model parameters fixed. For the diffusion-based SE, we set the number of reverse steps to N=25N{=}25.
@@ 4.3 Evaluation metrics
We evaluate adversarial attacks along two axes: attack success - how well the output signal $\hat{S}^{Y^{\mathrm{user}}+\delta}$ matches the target speech $S^{\mathrm{attacker}}$; and perturbation impact - how audible the perturbation $\delta$ is within the original mixture $Y^{\mathrm{user}}$.
Attack success (AS). We report WER, POLQA, ESTOI between $\hat{S}^{Y^{\mathrm{user}}+\delta}$ and $S^{\mathrm{attacker}}$, and DistillMOS [14] for $\hat{S}^{Y^{\mathrm{user}}+\delta}$. For WER, we transcribe with five Whisper [15] variants (Large, Large-v1, Large-v2, Large-v3, Large-v3-turbo) and average the resulting WERs for stability.
Perturbation impact (PI). To quantify how much the input mixture is altered, we use POLQA and ESTOI as a distance metric between $(Y^{\mathrm{user}}+\delta)$ and $Y^{\mathrm{user}}$.
@@
Here $\|\cdot\|_{2}$ denotes the Euclidean norm. A higher SNR indicates that the adversarial attack is of lower power.
@@
Table 2: Ablation of diffusion SE on the same 100 pairs. Unless specified, the sampler uses N=25N{=}25 reverse steps and stochastic sampling; Fixed-noise denotes a frozen noise path (Sec.  2.2). The baseline model is trained with $\sigma_{\mathrm{max}}{=}0.5$; rows that list $\sigma_{\mathrm{max}}$ are separately trained models.
@@ 5 Results and Analysis
Table 1 summarizes targeted attacks on 100 EARS-WHAM-v2 pairs. Below we highlight the main results.
@@
(1) Fixed energy ($\varepsilon{=}10$), sweep $\lambda$: with the energy budget fixed, SNR and perturbation impact metrics stay within a narrow band (SNR $\approx 8-13\ {dB}$; PI-POLQA $\approx 2.55-3.19$; PI-ESTOI $\approx 0.65-0.72$), while attack success shifts significantly - e.g., WER $\approx 0.20\to 1.37$, AS–ESTOI $\approx 0.77\to 0.22$, AS–POLQA $\approx 2.62\to 1.07$ (DistillMOS $\approx 3.27\to 1.72$) as $\lambda$ tightens ($40\to 0$).
(2) Fixed masking ($\lambda{=}20$), sweep $\varepsilon$: enlarging $\varepsilon$ improves alignment (WER $1.17\to 0.23$ from $\varepsilon{=}3\to\infty$) while increasing audibility (PI-POLQA $4.59\to 2.40$; SNR $32.46\to 7.72$ dB). Practically, moderate settings (e.g., $\lambda{=}20$, $\varepsilon\in\{10,15\}$) sit in a balanced region (WER $\approx 0.20-0.17$, PI-POLQA $\approx 3.19-2.26$, SNR $\approx 12.88-5.64$ dB).
CRM (predictive, mask). For identical $(\lambda,\varepsilon)$, perturbation impact matches Direct Map closely (e.g., at $\lambda{=}20,\varepsilon{=}10$: PI-POLQA $3.14$ vs. $3.19$; SNR $12.90$ vs. $12.88$ dB). Attack success is consistently, but only slightly lower (e.g., DistillMOS $2.54$ vs. $2.19$, AS-POLQA $1.81$ vs. $1.57$, WER $0.23$ vs. $0.20$ at the same setting). In short, the masking approach with similar perturbation impact metrics produces a worse attack.
@@ 6 Conclusions
In this paper, we show that modern speech enhancement systems exhibit some vulnerability to adversarial attacks. That is, by adding adversarial noise, an attacker can trick a speech enhancement system to output a completely different signal than intended by the user. To that end, we build white-box attacks with psychoacoustic masking and an $\ell_{2}$ budget for the adversarial attack. We test predictive (direct map, complex mask) and diffusion SE on EARS-WHAM-v2 using success and audibility metrics.
We show that predictive SE is comparably easily attacked with a relatively small $\ell_{2}$ budget for the adversarial noise. The complex mask variant behaves similarly to direct mapping but is somewhat less vulnerable. Diffusion SE is more resistant to targeted manipulation: even with a larger $\ell_{2}$ budget for the adversarial noise, the output aligns less well with the attacker’s target signal and remains more audible than predictive baselines. We show that stochastic sampling in the reverse diffusion process increases robustness to attacks even further.